OpenStudy (anonymous):

A=1/2aP What is the value of P if A= 40ft^2 nd a=8 ft?

OpenStudy (campbell_st):

if you substitute the value you have been given it becomes \[40 = \frac{1}{2} \times 8 \times P\] so you should be able to solve fo P now.